# L4: Low-Latency and Load-Balanced LLM Serving via Length-Aware Scheduling

### v1

@@ 6.5 Design Choice Validation and Ablation Study
Length-aware vs. naïve stage partitioning. We compare L4 with two naïve layouts: a chain layout, which assigns one instance per pipeline stage, and a no-pipeline layout, which coalesces all instances into a single stage. As shown in Figure˜14, the absence of length-based partitioning causes the no-pipeline design to perform the worst, while the chain layout underperforms L4 due to its overly long pipeline, which incurs higher migration overhead and complicates load balancing across stages. Consequently, L4 achieves 30% lower latency and 7.1% higher throughput compared to the chain layout.
(a) Normalized latency
(b) System throughput
Adaptive vs. naïve range-refinement policies. We compare L4 against two naïve range-refinement policies: a quantity-based refinement that balances the number of requests per stage, and a memory-based refinement that equalizes per-stage memory usage. Both policies enforce increasing request lengths across stages, as in L4. However, the quantity-based approach yields severe workload imbalance and thus the poorest performance, while the memory-based scheme fails to account for request heterogeneity as in Section˜4.3. As Figure˜15 illustrates, L4 outperforms the memory-based policy by 21% lower latency and 12% higher throughput.
@@ 7 Related Work
Attention backends. Modern self-attention kernels such as FlashAttention [dao2022flashattention] and FlashInfer [ye2025flashinfer] improve memory efficiency via I/O-aware designs and paged KV caches [kwon2023vllm]. Transformer libraries [fasttransformer, xFormers2022] and low-level frameworks like Triton [tillet2019triton], CUTLASS [cutlass], and CUDA [guide2020cuda] provide flexible platforms for scalable attention. L4 complements these by reducing request-level length heterogeneity to unlock the efficiency of the attention backend.
Intra-instance scheduling. vLLM [kwon2023vllm] and SGLang [sglang] dynamically form iteration batches using FCFS or length-aware policies. Prior work [zheng2023response] reduced output-length heterogeneity in traditional one-shot batching through output-length prediction, thereby reducing hardware idle time. However, this batching paradigm has been superseded by the more efficient continuous batching [yu2022orca] that underpins L4. FastServe [wu2023fast], SSJF [qiu2024efficient], and similar works [fu2024efficient] prioritize short-output requests to improve latency. Yet, this comes at the cost of throughput: effective prioritization often relies on small batch sizes (e.g., $\leq 10$ in SSJF) and aggressive preemption, leading to high swapping overhead [wu2023fast] that hinders system efficiency. L4 avoids this trade-off by distributing length-homogeneous batches across its instances.
Inter-instance scheduling and distributed inference. Multi-instance systems [sun2024llumnix, dynamo, team2025aibrix] like Llumnix [sun2024llumnix] employ load- and cache-aware heuristics to balance traffic among instances. Others [zhong2024distserve, zhu2025megascale, step3] decouple inference into phases (e.g., prefill vs. decode) and assign these phases to specialized instance pools or heterogeneous hardware. L4 is compatible with these approaches and enhances them via length-based instance partitioning.
Request migration in LLM serving. ServerlessLLM [fu2024serverlessllm] and Llumnix [sun2024llumnix] enable live request migration for stall-free handoff. L4 builds on this by integrating migration into the decentralized bid-ask mechanism for balancing length-specialized instances.
@@ 8 Discussion
Scope and limitations. L4 is most effective when ample GPU memory enables large-batch processing, shifting the primary latency bottleneck from parameter access to attention computation and KV cache intensity. Gains scale with available memory and parallelism. When the parameter size is large relative to the total GPU memory (e.g., occupying 80% of the available capacity) and the remaining KV cache memory cannot support large batches, the realized benefits are naturally constrained, yet the method remains effective. Workloads with highly uniform lengths inherently present less heterogeneity, where L4 still maintains correct scheduling and avoids overhead.
Orthogonality. L4 is independent of any particular inference engine. While our prototype inherits vLLM’s single-instance performance characteristics, the scheduling layer itself is complementary to kernel-level and intra-instance optimizations; faster execution engines would only amplify the benefits of global coordination.
Scalability. L4 combines offline bucket planning with lightweight online bid-ask rescheduling, keeping planning overhead low and KV migration infrequent. KV migration is efficient (Section˜5) and rarely impacts performance under realistic bandwidth; even under stress, L4 degrades gracefully.
Heterogeneous instances. While our evaluation employs homogeneous instance setups to align with standard production practices for simplified management, L4’s mechanisms—including pipeline planning, boundary adjustment, and bid-ask scheduling—natively support heterogeneous configurations. However, fully exploiting this flexibility presents a significant challenge: modern inference engines (e.g., vLLM and SGLang) expose hundreds of configurable parameters (e.g., tensor, pipeline, and expert parallelism degrees). Identifying the optimal configuration for each pipeline stage involves a high-dimensional search with prohibitive profiling costs, which we leave for future exploration.
@@ 9 Conclusion
By elevating length heterogeneity from a local batching issue to a cluster-level scheduling problem, L4 enables effective request regrouping across identical replicas, overcoming fundamental limits of per-instance schedulers. Through dynamic boundaries and decentralized rescheduling, L4 achieves substantial utilization and throughput gains, advancing scalable and efficient LLM serving.

### v2

@@ 6.5 Design Choice Validation and Ablation Study
Length-aware vs. naïve stage partitioning. We compare CascadeInfer with two naïve layouts: a chain layout, which assigns one instance per pipeline stage, and a no-pipeline layout, which coalesces all instances into a single stage. As shown in Figure˜14, the absence of length-based partitioning causes the no-pipeline design to perform the worst, while the chain layout underperforms CascadeInfer due to its overly long pipeline, which incurs higher migration overhead and complicates load balancing across stages. Consequently, CascadeInfer achieves 30% lower latency and 7.1% higher throughput compared to the chain layout.
Adaptive vs. naïve range-refinement policies. We compare CascadeInfer against two naïve range-refinement policies: a quantity-based refinement that balances the number of requests per stage, and a memory-based refinement that equalizes per-stage memory usage. Both policies enforce increasing request lengths across stages, as in CascadeInfer. However, the quantity-based approach yields severe workload imbalance and thus the poorest performance, while the memory-based scheme fails to account for request heterogeneity as in Section˜4.3. As Figure˜15 illustrates, CascadeInfer outperforms the memory-based policy by 21% lower latency and 12% higher throughput.
@@ 7 Related Work
Attention backends. Modern self-attention kernels such as FlashAttention [dao2022flashattention] and FlashInfer [ye2025flashinfer] improve memory efficiency via I/O-aware designs and paged KV caches [kwon2023vllm]. Transformer libraries [fasttransformer, xFormers2022] and low-level frameworks like Triton [tillet2019triton], CUTLASS [cutlass], and CUDA [guide2020cuda] provide flexible platforms for scalable attention. CascadeInfer complements these by reducing request-level length heterogeneity to unlock the efficiency of the attention backend.
Intra-instance scheduling. vLLM [kwon2023vllm] and SGLang [sglang] dynamically form iteration batches using FCFS or length-aware policies. Prior work [zheng2023response] reduced output-length heterogeneity in traditional one-shot batching through output-length prediction, thereby reducing hardware idle time. However, this batching paradigm has been superseded by the more efficient continuous batching [yu2022orca] that underpins CascadeInfer. FastServe [wu2023fast], SSJF [qiu2024efficient], and similar works [fu2024efficient] prioritize short-output requests to improve latency. Yet, this comes at the cost of throughput: effective prioritization often relies on small batch sizes (e.g., $\leq 10$ in SSJF) and aggressive preemption, leading to high swapping overhead [wu2023fast] that hinders system efficiency. CascadeInfer avoids this trade-off by distributing length-homogeneous batches across its instances.
Inter-instance scheduling and distributed inference. Multi-instance systems [sun2024llumnix, dynamo, team2025aibrix] like Llumnix [sun2024llumnix] employ load- and cache-aware heuristics to balance traffic among instances. Others [zhong2024distserve, zhu2025megascale, step3] decouple inference into phases (e.g., prefill vs. decode) and assign these phases to specialized instance pools or heterogeneous hardware. CascadeInfer is compatible with these approaches and enhances them via length-based instance partitioning.
Request migration in LLM serving. ServerlessLLM [fu2024serverlessllm] and Llumnix [sun2024llumnix] enable live request migration for stall-free handoff. CascadeInfer builds on this by integrating migration into the decentralized bid-ask mechanism for balancing length-specialized instances.
@@ 8 Discussion
Scope and limitations. CascadeInfer is most effective when ample GPU memory enables large-batch processing, shifting the primary latency bottleneck from parameter access to attention computation and KV cache intensity. Gains scale with available memory and parallelism. When the parameter size is large relative to the total GPU memory (e.g., occupying 80% of the available capacity) and the remaining KV cache memory cannot support large batches, the realized benefits are naturally constrained, yet the method remains effective. Workloads with highly uniform lengths inherently present less heterogeneity, where CascadeInfer still maintains correct scheduling and avoids overhead.
Orthogonality. CascadeInfer is independent of any particular inference engine. While our prototype inherits vLLM’s single-instance performance characteristics, the scheduling layer itself is complementary to kernel-level and intra-instance optimizations; faster execution engines would only amplify the benefits of global coordination.
Scalability. CascadeInfer combines offline bucket planning with lightweight online bid-ask rescheduling, keeping planning overhead low and KV migration infrequent. KV migration is efficient (Section˜5) and rarely impacts performance under realistic bandwidth; even under stress, CascadeInfer degrades gracefully.
Heterogeneous instances. While our evaluation employs homogeneous instance setups to align with standard production practices for simplified management, CascadeInfer’s mechanisms—including pipeline planning, boundary adjustment, and bid-ask scheduling—natively support heterogeneous configurations. However, fully exploiting this flexibility presents a significant challenge: modern inference engines (e.g., vLLM and SGLang) expose hundreds of configurable parameters (e.g., tensor, pipeline, and expert parallelism degrees). Identifying the optimal configuration for each pipeline stage involves a high-dimensional search with prohibitive profiling costs, which we leave for future exploration.
@@ 9 Conclusion
By elevating length heterogeneity from a local batching issue to a cluster-level scheduling problem, CascadeInfer enables effective request regrouping across identical replicas, overcoming fundamental limits of per-instance schedulers. Through dynamic boundaries and decentralized rescheduling, CascadeInfer achieves substantial utilization and throughput gains, advancing scalable and efficient LLM serving.